OpenStudy (anonymous):

Which arrangement shows 1.1 , 1 and 1/9 , 1.09 , and 9/8 in order from least to greatest? A. 1 and 1/9, 1.09, 9/8, 1.1 B. 1 and 1/9, 1.09, 1.1, 9/8 C. 1.09, 1 and 1/9, 1.1, 9/8 D. 1.09, 1.1, 1 and 1/9, 9/8

OpenStudy (timaashorty):

Change all intp a decimal \[\frac{ 1 }{ 9} = 0.11\] \[\frac{ 1 }{ 09 } = 0.1\] \[\frac{ 9 }{ 8 }= 1.13\] So we have 1.1, 1, 0.1, 0.11 , 1.13 in order they'll be: 0.1, 0.11, 1, 1.1, 1.13

OpenStudy (timaashorty):

Okay this is confusing. are you sure you wrote the answers correctly? cuz 1/09 = 1.9

OpenStudy (timaashorty):

OOPS i made a mistake

OpenStudy (timaashorty):

i thought it said 1/9

OpenStudy (anonymous):

Oh sorry for confusing you.

OpenStudy (timaashorty):

Not it's my fault anyways we know It's not A for sure because 9/8 is the greatest

OpenStudy (timaashorty):

B. is the correct answer. Sorry for the late response but I just saw the 'and'

OpenStudy (anonymous):

oh its okay thank you

OpenStudy (timaashorty):

You're welcome [:

OpenStudy (timaashorty):

If you need further help; just tag me

OpenStudy (anonymous):

Okay thank you.